Circular No. 7452

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports observations of twelve more Kreutz sungrazing comets detected with the SOHO/LASCO C2 coronagraph. The 1998 and 1999 comets were found by M. Oates from archival data in the SOHO website. The 2000 comets were found in current data in the website by M. Meyer (C/2000 M3), by Meyer and M. Boschat (C/2000 M4) and by A. Mimeev and P. Shkreby (C/2000 M5). Positions have been measured by D. Hammer (in collaboration with Biesecker in the case of the very faint and diffuse object C/2000 M3) and reduced by B. G. Marsden. Discovery positions follow. The complete measurements and orbital elements are given on MPEC 2000-N26.

1998 UT  α_{2000}  δ_{2000}  Comet
June 6.644  4^{h}54.^{m}7  +20^{o}50′  C/1998 L7
8.334  5 02.0  +21 02  C/1998 L8
12.226  5 18.7  +21 21  C/1998 L9
19.414  5 46.8  +21 44  C/1998 M8
21.622  5 55.3  +21 45  C/1998 M9
23.644  6 02.9  +21 59  C/1998 M10
1999 UT  α_{2000}  δ_{2000}  Comet
June 1.299  4^{h}34.^{m}6  +20^{o}18′  C/1999 L6
3.129  4 40.5  +20 25  C/1999 L7
6.310  4 54.2  +20 53  C/1999 L8
2000 UT  α_{2000}  δ_{2000}  Comet
June 19.347  5^{h}49.^{m}0  +21^{o}41′  C/2000 M3
22.121  5 59.9  +21 38  C/2000 M4
24.132  6 07.0  +22 00  C/2000 M5

SUPERNOVA 2000ck IN IC 4355 A. V. Filippenko, R. Chornock and M. Modjaz, University of California

at Berkeley, report that inspection of a CCD spectrum (range 320–1000 nm) of supernova 2000ck obtained on July 6 with the Shane 3-m reflector at the Lick Observatory confirms that the object is a type II supernova (cf. IAUC 7434). Weak (though distinct) broad Hα emission has appeared, but there is essentially no absorption component to the profile. This behavior is similar to that of the type II-L supernovae 1979C and 1980K (see Filippenko 1997, Ann. Rev. A. Ap. 35, 309 for a relevant discussion).

2000 July 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden